OpenStudy (anonymous):

The value of 5(1 + 3)^2 is 400. I did 1+3 squared so 8 x 5 = 40?

OpenStudy (anonymous):

5(1+3)^2 does not equal 400, but (5(1+3))^2 does (5(1+3))^2=400 (5(4))^2=400 20^2=400 400=400

OpenStudy (dan815):

alexis..

OpenStudy (anonymous):

yes? @dan815